OpenStudy (anonymous):

Can somebody Help Me Please! ? Mason delivers newspapers to 400 homes in his community. He can deliver an average of 3 newspapers a minute. Write an expression that represents how many newspapers he has left to deliver after m minutes.

OpenStudy (anonymous):

i think, newspaper left after m minute is=400-3m also with condition \[400-3m \ge 0\]

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

that was the answer. expression for newspaper left after m minute is =400-3m

OpenStudy (anonymous):

Thanks Bro